OpenStudy (anonymous):

Solve the equation |27-75x|=-1000 .

OpenStudy (anonymous):

i think i just have to solve each the negative and positive one right?

OpenStudy (anonymous):

yes/

OpenStudy (jamesj):

Is there ANY number y such that |y| = -1? Or such that |y| = -1000?

OpenStudy (anonymous):

thanks!

OpenStudy (jamesj):

No; look at what I'm saying.

OpenStudy (anonymous):

what do you mean?

OpenStudy (jamesj):

Your problem has this expression: |27-75x|=-1000 But can there be any number y such that |y| = -1000? No, there can't because |y| is always positive or zero. |y| can never have a negative value. So |27-75x|=-1000 is meaningless.

OpenStudy (jamesj):

Hence the equation |27-75x|=-1000 has NO solutions.

OpenStudy (anonymous):

when i was told to find this answer i was told to get 2 answers one negative and one positive

OpenStudy (turingtest):

James is right, ignore FoolforMath.

OpenStudy (jamesj):

That works for | something | = a positive number or zero but if | something | = negative number, then the expression is inconsistent/meaningless/invalid and has no solutions.

OpenStudy (anonymous):

@TuringTest: I assumed it's a inequality as the OP is mentioning for solving from both sides. However,things can be made pretty clear only if you define abs value, Absolute value of a number is its distance from zero on the number line. The absolute value of a number n is denoted by, |n|. So,asking for solving |27-75x|=-1000 . is meaningless and insane and stupid ... so I OP meant an inequality instead of equality.

OpenStudy (anonymous):

OH i never knew that... so since 1000 is negative that makes the equation no solution?

OpenStudy (jamesj):

If you're not convinced, let me ask you: what number y has the property that | y | = -1000? Well y = 1000 doesn't work because | 1000 | = 1000 and y = -1000 doesn't work either because | -1000 | = 1000. So yes, the equation as posed has no solutions.

OpenStudy (anonymous):

@MoreForJoe: First learn the definition and then solve the problem :)

OpenStudy (anonymous):

i just never knew about the negative part. our professor just threw us a curve-ball

OpenStudy (anonymous):

and by number JamesJ means real numbers only,absolute value of complex numbers is a simmilar concept but computation is a bit different ...:)

OpenStudy (anonymous):

@MoreForJoe: Let me remind you the definition again ... *Absolute value of a number is its distance from zero on the number line.* Distance being a scalar quantity, how can it be negative ? ;)

OpenStudy (turingtest):

like James said\[\left| y \right|=-x\]is IMPOSSIBLE because of the concept of the absolute value. @FoolforMath, if this was an inequality it would still have no solutions\[\left| y \right|\le-x\]cannot be broken into\[-x \le y \le x\]for the same reasons as the equality cannot. think about why

OpenStudy (anonymous):

ahhhh ok

OpenStudy (anonymous):

JamesJ didn't mentioned the definition anywhere and mostly explained on examples ... oh yes I can see that how can a distance be less than or equal to a negative number ? :O but well \[ \left| y \right|\ge-x \] would have and so do \[ \left| y \right|\le-x \] with x=-z,where \[ x,y,z \in \mathbb{R} \] ;)

OpenStudy (turingtest):

That's a little far-fetched reasoning don't you think?

OpenStudy (jamesj):

Fine. By definition \[ |x| = \left\{ \begin{matrix}x & \text{ if x > 0} \\ 0 & \text{ if x = 0} \\ -x & \text{ if x < 0}\end{matrix} \right. \] Hence if x > 0, then |x| = x > 0 if x = 0, then |x| = 0 if x < 0, then |x| = -x > 0 Therefore \[|x| \geq 0\] for all values of x.

OpenStudy (anonymous):

@TuringTest:Well,on second thought,may be not as you asked me that question and want me to think over it.

OpenStudy (jamesj):

Therefore, for any negative number a, there are no real numbers x such that \[ |x| = a < 0 \]

OpenStudy (anonymous):

what about something like that?

OpenStudy (anonymous):

Let f(x)=|2x 10| and g(x)=|2x 24| . Find all values of x for which f(x)=g(x) .

OpenStudy (anonymous):

if they are equal to each other?

OpenStudy (anonymous):

what is |2x 10| ? Are you missing any operator ?

OpenStudy (anonymous):

oh! oops Let f(x)=|2x+10| and g(x)=|2x-24| . Find all values of x for which f(x)=g(x) .

OpenStudy (anonymous):

May be you could post that as new question ...

OpenStudy (anonymous):

i tried no one replied but ill post it again

OpenStudy (anonymous):

If you don't get answers from this site .. post it here : math.stackexchange.com